Bauer Intelligence (BI) is a publicly traded company with a current share price of $20 per share. BI has 30 million shares outstanding, $80 million in debt, and $12 million in cash. BI plans to pay $1.50 per share in dividends in the coming year and the dividends are expected to grow by 4% per year in the future. BI’s long-term debt consists of bonds issued with a face value of $80 million with 10 years to maturity with annual coupon rate of 10% (APR). The long-term bonds are currently trading at par value. The beta of BI is 1.30, the risk-free rate is 3%, and the required return on the market portfolio is 8%. The corporate tax rate is 30%.

  1. What is the BI’s cost of equity based on dividend discount model (DDM)?  (2 marks)
  1. What is the BI’s cost of equity based on capital asset pricing model (CAPM)?  (2 marks)
  2. What would the growth rate in DDM has to be in order to reach the same consensus as CAPM? (2 marks)
  3. What is the BI’s after-tax cost of debt?  (2 marks)

  1. Using the cost of equity obtained using CAPM in part (b) above, calculate BI’s weighted average cost of capital (WACC)?                                                                           (3 marks)
Answer #1

a) As per dividend discount model , cost of equity = Expected dividend/Price per share + growth rate

Growth rate = 4%

Price per share = 20$

Expected discount = 1.50 $

Thus Cost of equity = 1.50/20 + 4%

= 0.075 + 0.04

= 11.5%

a) As per CAPM , cost of equity = Risk free rate of return + Beta(Market return - Risk free rate of return)

= 3% + 1.3(8%-3%)

=3% + 1.3(5%)

= 3%+ 6.5%

= 9.5%

b) For  same consensus as CAPM, Growth rate should be

Cost of equity as per CAPM = cost of equity ss per dividend discount model

= 9.5% = Expected dividend/Price per share + growth rate

= 9.5% = 1.5/20 + Growth rate

0.095 = 0.075 + Growth rate

Growth rate = 0.020

i.e. 2%

c) After tax cost of debt = Annual coupon rate (1-tax rate)

Since bonds are trading at par, Coupon rate = YTM = 10%

Thus After tax cost of debt = 10%(1-tax rate)

= 10%(1-30%)

= 10%(0.7)

= 7%

d) Statement showng WACC

Particulars Amount Weight Cost of capital WACC
(in million) a b c =axb
Equity 600.00 88% 9.500% 8.38%
Debt 80.00 12% 7.000% 0.82%
WACC 680.00 9.21%

Thus WACC = 9.21%